The volume of a gas is 275 mL at 740 torrs and 25°C. If the pressure is held constant and the temperature is raised to 100°C, what will be it's new volume


1
Expert's answer
2021-11-08T08:12:14-0500

"V_1 = 275 \;mL \\

p_1 = p_2 = 740 \; torrs \\

T_1 = 25+273 = 298 \;K \\

T_2 = 100 + 273 = 373 \;K \\

\frac{p_1V_1}{T_1} = \frac{p_2V_2}{T_2} \\

p = constant \\

\frac{V_1}{T_1} = \frac{V_2}{T_2} \\

V_2 = \frac{V_1T_2}{T_1} \\

V_2 = \frac{275 \times 373}{298} = 344.2 \;mL"

Answer: 344 mL